OpenStudy (anonymous):

Which of the following cereals is the best buy? A. 28 oz. of cereal for $4.29 B. 32 oz. of cereal for $4.59 C. 36 oz. of cereal for $4.79 D. 24 oz. of cereal for $3.59

OpenStudy (anonymous):

i would go with c

OpenStudy (black_ninja123):

Yea i would also go for C just divide the numbers

OpenStudy (texaschic101):

A. 4.29/28 = 0.1532 rounds to 0.15 per ounce B. 4.59/32 = 0.1434 rounds to 0.14 per ounce C. 4.79/36 = 0.1330 rounds to 0.13 per ounce D. 3.59/24 = 0.1495 rounds to 0.15 per ounce best buy (cheapest one per ounce) is C

OpenStudy (anonymous):

Preciate it! @texaschic101 @Black_ninja123 @DragonsGirl